OpenStudy (anonymous):

find the zeros of the polynomial function f(x)=x^2-12x+20

OpenStudy (anonymous):

this one factors as \[(x-2)(x-10)=0\] solve that one

OpenStudy (anonymous):

answers a) 2, 10 b)-4, -5 c) -2,-10 d) 4,5

OpenStudy (anonymous):

0

OpenStudy (anonymous):

c?

OpenStudy (jim766):

break it apart and make each = 0 x - 2 = 0 or x -10 = 0 now solve for x

OpenStudy (jim766):

add 2 to both sides x = 2

OpenStudy (jim766):

add 10 to both sides x = 10

OpenStudy (anonymous):

so 2,10

OpenStudy (jim766):

right